Cheyenne Corp. had record sales in 2020. It began 2020 with an Accounts Receivable balance of $455,000 and an Allowance for Doubtful Accounts of $32,000. Cheyenne recognized credit sales during the year of $6,650,000 and made monthly adjusting entries equal to 0.5% of each month’s credit sales to recognize bad debt expense. Also during the year, the company wrote off $36,900 of accounts that were deemed to be uncollectible, although one customer whose $5,400 account had been written off surprised management by paying the amount in full in late September. Including this surprise receipt, $6,536,000 cash was collected on account in 2020.

To assess the reasonableness of the allowance for doubtful accounts, the controller prepared the following aged listing of the receivables at December 31, 2020:
Days Account Outstanding Amount Probability of Collection
Less than 16 days $261,900 97%
Between 16 and 30 days 115,000 90%
Between 31 and 45 days 78,000 82%
Between 46 and 60 days 38,100 73%
Between 61 and 75 days 20,400 51%
Over 75 days 24,100 0%
$537,500

Reconcile the 2020 opening balance in Accounts Receivable to the $537,500 ending balance on the controller’s aged listing.

Prepare the adjusting entry to bring the Allowance for Doubtful Accounts to its proper balance at year end.

Show how accounts receivable would be presented on the December 31, 2020 statement of financial position.

What is the dollar effect of the year-end bad debt adjustment on the before-tax income?

The year-end bad debt adjustment would

increase or decrease

before-tax income $=

.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Thanks!
Cheyenne Corp.
Reconciliation of Accounts Receivable Amount $
Opening Balance           455,000.00
Add: Credit sales during the year        6,650,000.00
Less: Bad debts written off              36,900.00
Add: Bad debts recovered                5,400.00
Less: cash collected        6,536,000.00
Closing Balance           537,500.00
Balance as given in question           537,500.00
Difference                            -  
Hence reconciled
Allowance for Doubtful Accounts
Opening Balance              32,000.00
Add: Allowance made during the year              33,250.00
(0.5% of monthly credit sales)
Less: Bad debts written off              36,900.00
Add: Bad debts recovered                5,400.00
Closing Balance (before adjustment)             33,750.00 E
Allowance for Doubtful Accounts that should be as on Dec 31, 2020.
A B C=1-B D=A*C
Days Account Outstanding Amount Probability of Collection Probability Uncollectible Allowance for Doubtful Accounts
Less than 16 days           261,900.00 97% 3%      7,857.00
Between 16 and 30 days           115,000.00 90% 10% 11,500.00
Between 31 and 45 days              78,000.00 82% 18% 14,040.00
Between 46 and 60 days              38,100.00 73% 27% 10,287.00
Between 61 and 75 days              20,400.00 51% 49%      9,996.00
Over 75 days              24,100.00 0% 100% 24,100.00
          537,500.00 77,780.00
Bad debt expense for Dec 31, 2020.
Allowance for Doubtful Accounts              77,780.00 See D
Already in the books              33,750.00 See E
Bad debt expense for Dec 31, 2020.             44,030.00 F=D-E
Adjusting Entry
Account Debit $ Credit $
Bad debt expense              44,030.00
Allowance for doubtful account      44,030.00
Balance sheet- Dec 31
Accounts Receivable (gross)           537,500.00
Less: Allowance for doubtful account              77,780.00
Accounts Receivable (net)           459,720.00
What is the dollar effect of the year-end bad debt adjustment on the before-tax income?
Answer $ 44,030.